OpenStudy (anonymous):

Solve for x: ax^2=bx

OpenStudy (anonymous):

Do you know how to isolate the variable?

OpenStudy (anonymous):

No, but I moved a to the other side, and now have x^2=bx/a.

OpenStudy (anonymous):

What do I do now?

OpenStudy (anonymous):

Because x is being multiplied by, you have to do the opposite on the right side

OpenStudy (anonymous):

You're on the right track

OpenStudy (anonymous):

Ahh, so when I do that, I get x=bx/ax. The x's cancel out, giving me x=b/a?

OpenStudy (anonymous):

correct, good job

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

Thank you for the medal